# RoadBench: A Vision-Language Foundation Model and Benchmark for Road Damage Understanding

### v1

@@ 5.2 Results
We compare RoadCLIP with a range of vision-only and multimodal models. As shown in Table 2, vision-only detectors such as YOLOv10 [60], PP-PicoDet [79], and RT-DETR [89] achieve moderate performance in Semantic Localization Accuracy (SLA), with TD-YOLOv10 reaching the highest score of 49.0%. However, these models lack the ability to leverage language supervision, limiting their performance in retrieval and zero-shot classification tasks. Multimodal vision-language models show stronger results across all metrics. GPT-4V [48] and DeepSeek-VL [43] achieve 74.5% and 72.9% in zero-shot accuracy (ZS Acc.), respectively, and perform competitively in image-text retrieval. In contrast, RoadCLIP achieves consistent improvements across all evaluation tasks. It obtains 78.6% in ZS Acc., outperforming GPT-4V and DeepSeek-VL by +4.1% and +5.7%, respectively. On retrieval, RoadCLIP achieves 58.4% Recall@1 and 84.6% Recall@10, establishing new state-of-the-art performance on RoadBench. In SLA, RoadCLIP surpasses GPT-4V by +6.5% and DeepSeek-VL by +8.9%, highlighting its advantage in spatial grounding. These results verify the effectiveness of our proposed Disease-aware Positional Encoding (DaPE) and domain-specific prior injection strategy. The improvements are achieved without relying on large-scale general-domain pretraining, but rather through targeted multimodal alignment tailored for the road damage domain. In summary, RoadCLIP provides a strong and reliable foundation for multimodal understanding in structured environments. We similarly provide comparative experiments with purely visual metrics, see supplementary materials for more details.
@@ 5.3 Evaluating the Effectiveness of DaPE
Method TD-RD[70] (ZS Acc. %) CNRDD[83] (Recall@1 %) CRDDC’22[3] (SLA %)
Score $\Delta$ vs. CLIP Rank Score $\Delta$ vs. CLIP Rank Score $\Delta$ vs. CLIP Rank
CLIP [ICML21] [50] 65.7 - 6 38.2 - 6 42.5 - 6
BLIP-2 [ICML23] [32] 67.4 +1.7 5 40.6 +2.4 5 44.9 +2.4 5
LLaVA [NeurIPS23] [40] 69.1 +3.4 4 41.7 +3.5 4 47.1 +4.6 2
GPT-4V [arXiv23] [48] 71.2 +5.5 2 44.9 +6.7 2 46.8 +4.3 3
DeepSeek-VL [arXiv24] [43] 70.5 +4.8 3 43.3 +5.1 3 45.2 +2.7 4
@@
5.4 Effectiveness of Multimodal Fusion
We conduct ablation experiments across single-modal and multimodal configurations to assess each modality’s contribution and the importance of domain-specific priors. The results are summarized in Table 5. RoadCLIP, which benefits from domain-specific prior injection, achieves state-of-the-art performance across all metrics. It outperforms the strongest alternative by +6.1% in ZS Acc., +8.6% in Recall@1, and +9.3% in SLA. These substantial gains underscore the importance of vision-language fusion for capturing detailed spatial semantics and improving cross-modal alignment.
Table 5: Ablation study on multimodal fusion effectiveness.
Modality Combination ZS Acc.(%) Recall@1(%) Recall@5(%) SLA(%)
Vision Only (Baseline CNN) 58.2 30.9 48.3 35.2
Vision Only (Baseline Transformer) 60.7 33.1 50.2 37.8
Text Only (Text Encoder) 48.6 - - -
Vision + Text (CLIP Original) 63.8 40.3 63.2 41.5
Vision + Text (BLIP-2 Fusion) 67.9 43.7 66.5 46.3
Vision + Text + Absolute Position Encoding 69.2 46.2 68.4 48.7
Vision + Text + Relative Position Encoding 72.5 49.8 72.9 52.6
Vision + Text + Disease Prior (RoadCLIP) 78.6 58.4 78.1 61.9
The experiments highlight three important findings: Single-modality approaches have clear limitations. Vision-only and text-only models consistently underperform, confirming that relying on a single modality is insufficient for accurately understanding road defects. Even with a Transformer backbone, the vision-only model achieves only 60.7% ZS Acc. and 33.1% Recall@1, while the text-only encoder fails to facilitate retrieval-based evaluation. These results reflect the inherent limitations and ambiguities when using either modality alone. Fusing visual and textual modalities yields significant improvements. CLIP and BLIP-2 substantially outperform unimodal baselines, demonstrating the effectiveness of multimodal alignment for road damage analysis. Specifically, BLIP-2 improves Recall@1 to 43.7% and SLA to 46.3%, validating the benefit of incorporating textual guidance into visual representation learning. Explicit positional encoding plays a crucial role. Relative positional encoding achieves notable performance gains, attaining 72.5% ZS Acc. and 49.8% Recall@1. This indicates that spatial cues—such as crack location—are essential for distinguishing fine-grained road defect categories and strengthening attention mechanisms.
@@ 5.5 Qualitative Analysis and Visualization
Figure 6: Explainability of RoadCLIP via heatmap++ [9]. RoadCLIP demonstrates superior attention localization on road damage regions compared to GPT-4o and DeepSeek-VL.
@@ 6 Conclusion
Acknowledgments
This manuscript was authored by Oak Ridge National Laboratory (ORNL), operated by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. Any subjective views or opinions expressed in this paper do not necessarily represent those of the U.S. Department of Energy or the United States Government.

### v3

@@ 5.2 Results
We compare RoadCLIP with a range of vision-only and multimodal models. As shown in Table 4.4, vision-only detectors such as YOLOv10 [wang2024yolov10], PP-PicoDet [yu2021pppicodetbetterrealtimeobject], and RT-DETR [Zhao_2024_CVPR] achieve moderate performance in Semantic Localization Accuracy (SLA), with TD-YOLOv10 reaching the highest score of 49.0%. However, these models lack the ability to leverage language supervision, limiting their performance in retrieval and zero-shot classification tasks. Multimodal vision-language models show stronger results across all metrics. GPT-4V [openai2023gpt4] and DeepSeek-VL [lu2024deepseekvl] achieve 74.5% and 72.9% in zero-shot accuracy (ZS Acc.), respectively, and perform competitively in image-text retrieval. In contrast, RoadCLIP achieves consistent improvements across all evaluation tasks. It obtains 78.6% in ZS Acc., outperforming GPT-4V and DeepSeek-VL by +4.1% and +5.7%, respectively. On retrieval, RoadCLIP achieves 58.4% Recall@1 and 84.6% Recall@10, establishing new state-of-the-art performance on RoadBench. In SLA, RoadCLIP surpasses GPT-4V by +6.5% and DeepSeek-VL by +8.9%, highlighting its advantage in spatial grounding. These results verify the effectiveness of our proposed Disease-aware Positional Encoding (DaPE) and domain-specific prior injection strategy. The improvements are achieved without relying on large-scale general-domain pretraining, but rather through targeted multimodal alignment tailored for the road damage domain. In summary, RoadCLIP provides a strong and reliable foundation for multimodal understanding in structured environments. We similarly provide comparative experiments with purely visual metrics, see supplementary materials for more details.
@@ 5.3 Evaluating the Effectiveness of DaPE
[…]
BLIP-2 [ICML23] [li2023blip] 67.4 +1.7 5 40.6 +2.4 5 44.9 +2.4 5
LLaVA [NeurIPS23] [liu2023llava] 69.1 +3.4 4 41.7 +3.5 4 47.1 +4.6 2
GPT-4V [arXiv23] [openai2023gpt4] 71.2 +5.5 2 44.9 +6.7 2 46.8 +4.3 3
DeepSeek-VL [arXiv24] [lu2024deepseekvl] 70.5 +4.8 3 43.3 +5.1 3 45.2 +2.7 4
@@ 5.4 Qualitative Analysis and Visualization
Figure 6: Explainability of RoadCLIP via heatmap++ [chefer2022attentive]. RoadCLIP demonstrates superior attention localization on road damage regions compared to GPT-4o and DeepSeek-VL.
@@ 6 Conclusion
Acknowledgment
This manuscript was co-authored by Oak Ridge National Laboratory (ORNL), operated by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. Any subjective views or opinions expressed in this paper do not necessarily represent those of the U.S. Department of Energy or the United States Government.